Let X and Y have joint probability distribution function f(x, y) = ( 2x+y /12) , (x, y) = (0, 1); (0, 2); (1, 2); (1, 3) 0, elsewhere

Find

(i) the covariance between X and Y.

(ii) the joint probability generating function of X and Y.


1
Expert's answer
2022-04-04T17:50:23-0400

"f\\left( x,y \\right) =\\left\\{ \\begin{array}{c}\t\\frac{1}{12},\\left( x,y \\right) =\\left( 0,1 \\right)\\\\\t\\frac{1}{6},\\left( x,y \\right) =\\left( 0,2 \\right)\\\\\t\\frac{1}{3},\\left( x,y \\right) =\\left( 1,2 \\right)\\\\\t\\frac{5}{12},\\left( x,y \\right) =\\left( 1,3 \\right)\\\\\\end{array} \\right. \\\\i:EX=0\\cdot \\frac{1}{12}+0\\cdot \\frac{1}{6}+1\\cdot \\frac{1}{3}+1\\cdot \\frac{5}{12}=\\frac{3}{4}\\\\EY=1\\cdot \\frac{1}{12}+2\\cdot \\frac{1}{6}+2\\cdot \\frac{1}{3}+3\\cdot \\frac{5}{12}=\\frac{7}{3}\\\\EXY=0\\cdot \\frac{1}{12}+0\\cdot \\frac{1}{6}+2\\cdot \\frac{1}{3}+3\\cdot \\frac{5}{12}=\\frac{23}{12}\\\\cov\\left( X,Y \\right) =\\frac{23}{12}-\\frac{3}{4}\\cdot \\frac{7}{3}=\\frac{1}{6}\\\\ii:\\\\M_{XY}\\left( t,s \\right) =Ee^{tX+sY}=e^s\\cdot \\frac{1}{12}+e^{2s}\\cdot \\frac{1}{6}+e^{t+2s}\\cdot \\frac{1}{3}+e^{t+3s}\\cdot \\frac{5}{12}"